# Unveiling Transformer Perception by Exploring Input Manifolds

### v1

@@ Abstract
This paper introduces a general method for the exploration of equivalence classes in the input space of Transformer models. The proposed approach is based on sound mathematical theory which describes the internal layers of a Transformer architecture as sequential deformations of the input manifold. Using eigendecomposition of the pullback of the distance metric defined on the output space through the Jacobian of the model, we are able to reconstruct equivalence classes in the input space and navigate across them. We illustrate how this method can be used as a powerful tool for investigating how a Transformer sees the input space, facilitating local and task-agnostic explainability in Computer Vision and Natural Language Processing tasks.
@@ 1 Introduction
In this paper, we propose a method for exploring the input space of Transformer models by identifying equivalence classes with respect to their predictions. We define an equivalence class of a Transformer model as the set of vectors in the embedding space whose outcomes under the Transformer process are the same. The study of the input manifold on which the inverse image of models lies provides insights for both explainability and sensitivity analyses. Existing methods aiming at the exploration of the input space of Deep Neural Networks and Transformers either rely on perturbations of input data using heuristic or gradient-based criteria [16, 22, 17, 14], or they analyze specific properties of the embedding space [5].
Our approach is based on sound mathematical theory which describes the internal layers of a Transformer architecture as sequential deformations of the input manifold. Using eigendecomposition of the pullback of the distance metric defined on the output space through the Jacobian of the model, we are able to reconstruct equivalence classes in the input space and navigate across them. In the XAI scenario, our framework can facilitate local and task-agnostic explainability methods applicable to Computer Vision (CV) and Natural Language Processing (NLP) tasks, among others.
In Section 2, we summarise the preliminaries of the mathematical foundations of our approach. In Section 3, we present our method for the exploration of equivalence classes in the input of the Transformer models. In Section 4, we perform a preliminary investigation of some applicability options of our method on textual and visual data. In Section 5, we discuss the relevant literature about embedding space exploration and feature importance. Finally, in Section 6, we give our concluding remarks [^1] The code to reproduce our experiments can be found in the Supplementary Materials..
@@ 2 Preliminaries
In this Section, we provide the theoretical foundation of the proposed approach, namely the Geometric Deep Learning framework based on Riemannian Geometry [2].
@@
[…]
We call $M_{0}$ the input manifold and $M_{n}$ the output manifold. All the other manifolds of the sequence are called representation manifolds. The maps $\Lambda_{i}$ are the layers of the neural network. We denote with $\mathcal{N}_{(i)}=\Lambda_{n}\circ\dots\circ\Lambda_{i}:M_{i}\to M_{n}$ the mapping from the $i$-th representation layer to the output layer.
@@
A map $\Lambda_{i}:M_{i-1}\rightarrow M_{i}$ is called a smooth layer if it is the restriction to $M_{i-1}$ of a function $\overline{\Lambda}^{(i)}(x):\mathbb{R}^{d_{i-1}}\rightarrow\mathbb{R}^{d_{i}}$ of the form
$\overline{\Lambda}^{(i)}_{\alpha}(x)=F^{(i)}_{\alpha}\left(\sum_{\beta}A_{\alpha\beta}^{(i)}x_{\beta}+b^{(i)}_{\alpha}\right)$ (2)
for $i=1,\cdots,n$, $x\in\mathbb{R}^{d_{i}}$, $b^{(i)}\in\mathbb{R}^{d_{i}}$ and $A^{(i)}\in\mathbb{R}^{d_{i}\times d_{i-1}}$, with $F^{(i)}:\mathbb{R}^{d_{i}}\rightarrow\mathbb{R}^{d_{i}}$ a diffeomorphism.
@@
Transformers implicitly apply for this framework, since their modules are smooth functions, such as fully connected layers, GeLU and sigmoid activations.
Our aim is to transport the geometric information on the data lying in the output manifold to the input manifold: this allows us to obtain insight on how the network "sees" the input space, how it manipulates it for reaching its final conclusion. For fulfilling this objective, we need several tools from differential geometry. The first key ingredient is the notion of singular Riemannian metric, which has the intuitive meaning of a degenerate scalar product which changes point to point.
@@
Without loss of generality, we can assume the following hypotheses on the sequence (1): i) The manifolds $M_{i}$ are open and path-connected sets of dimension $\dim M_{i}=d_{i}$. ii) The maps $\Lambda_{i}$ are $\mathcal{C}^{1}$ submersions. iii) $\Lambda_{i}(M_{i-1})=M_{i}$ for every $i=1,\cdots,n$. iv) The manifold $M_{n}$ is equipped with the structure of Riemannian manifold, with metric $g^{(n)}$. Definition 3 naturally leads to the definition of the pseudolength and of energy of a curve.
@@
Let $\gamma:[a,b]\rightarrow\mathbb{R}^{n}$ a curve defined on the interval $[a,b]\subset\mathbb{R}$ and $\|v\|_{p}=\sqrt{g_{p}(v,v)}$ the pseudo–norm induced by the pseudo–metric $g_{p}$ at point $p$. Then the pseudolength of $\gamma$ and its energy are defined as
[…]
@@
The second crucial tool is the notion of pullback of a function. Let $f$ be a function from $\mathbb{R}^{p}$ to $\mathbb{R}^{q}$, and fix the coordinate systems $x=(x_{1},\dots,x_{p})$ and $y=(y_{1},\dots,y_{q})$ on $\mathbb{R}^{p}$ and on $\mathbb{R}^{q}$, respectively. Moreover, we endow $\mathbb{R}^{q}$ with the standard Euclidean metric $g$, whose associated matrix is the identity. The space $\mathbb{R}^{p}$ can be equipped with the pullback metric $f^{*}g$ whose representation matrix reads as
@@
Hereafter, we consider in (1) the case $M_{n}=\mathbb{R}^{q}$, equipped with the trivial metric $g^{(n)}=I_{q}$, i.e., the identity. Each manifold $M_{i}$ of the sequence (1) is equipped with a Riemannian singular metric, denoted with $g^{(i)}$, obtained via the pullback of $\mathcal{N}_{(i)}$. The pseudolength of a curve $\gamma$ on the $i$-th manifold, namely $Pl_{i}(\gamma)$, is computed via the relative metric $g^{(i)}$ via (3).
@@ 2.1 General results
Let $\gamma:[0,1]\rightarrow M_{i}$ be a piecewise $\mathcal{C}^{1}$ curve. Let $k\in\{i,i+1,\cdots,n\}$ and consider the curve $\gamma_{k}=\Lambda_{k}\circ\cdots\circ\Lambda_{i}\circ\gamma$ on $M_{k}$. Then $Pl_{i}(\gamma)=Pl_{k}(\gamma_{k})$.
@@
The following corollary contains the natural consequences of the previous result; the second point of the claim below is the counterpart of 2.
@@
Under the hypothesis of 3, one has that ${M_{i}}/{\sim_{i}}={M_{i}}/{\sim_{\mathcal{N}_{i+1}}}$. Moreover, if two points $p,q\in M_{i}$ are connected by a $\mathcal{C}^{1}$ curve $\gamma:[0,1]\rightarrow M_{i}$ satisfying $\mathcal{N}_{i}(p)=\mathcal{N}_{i}\circ\gamma(s)$ for every $s\in[0,1]$, then they lie in the same class of equivalence.
Making use of the Godement’s criterion, we are now able to prove that the set $M_{i}/\sim_{i}$ is a smooth manifold, together with its dimension.
@@
This last achievement provides practical insights about the projection $\pi_{i}$ on the quotient space, that consists the building block of the algorithms used for recovering and exploring the equivalence classes of a neural network.
@@ 3 Methodology
The results depicted in Section 2.1 provide powerful tools for investigating how a neural network sees the input space starting from a point $x$. In particular we point out the following remarks: i) If two points $x,y$ belonging to the input manifold $M_{0}$ are are such that $x\sim_{0}y$, then $\mathcal{N}(x)=\mathcal{N}(y)$; ii) given a point $p\in M_{n}$, the counterimage $\mathcal{N}^{-1}(p)$ is a smooth manifold, whose connected components are classes of equivalences in $M_{0}$ with respect to $\sim_{0}$. A necessary condition for two points $x,y\in M_{0}$ to be in the same class of equivalence is that $\mathcal{N}(x)=\mathcal{N}(y)$; iii) any class of equivalence $[x]$, $x\in M_{0}$, is a maximal integral submanifold of $\mathcal{V}M_{0}$. The above observations directly provide with a strategy to build up the equivalence class of an input point $x\in M_{0}$. 5 tells us that $\mathcal{V}M_{0}$ is an integrable distribution, with dimension equal to the dimension of the kernel of $g^{(0)}$: we can hence find $dim(Ker(g^{(0)}))$ vector fields which are a base for the tangent space of $M_{0}$. This means that we can compute the eigenvalue decomposition of $g^{(0)}_{x}$ and consider the $L$ linearly independent eigenvectors, namely $\{v_{l}\}_{l=1,\dots,L}$, associated to the null eigenvalue: these eigenvectors depend smoothly on the point, a fact that is not trivial when the matrix associated to the metric depends on several parameters [15]. We can build then all the null curves by randomly selecting one eigenvector $\tilde{v}\in\{v_{l}\}$ and then reconstruct the curve along the direction $\tilde{v}$ from the starting point $x$. From a practical point of view, one is led to solve the Cauchy problem, a first order differential equation, with $\dot{\gamma}=\tilde{v}$ and initial condition $\gamma(0)=x$.
@@ 3.1 Input Space Exploration
This whole procedure is coded in the Singular Metric Equivalence Class (SiMEC) and the Singular Metric Exploration (SiMExp) algorithms, whose general schemes are depicted in Algorithms 1 and 2. SiMEC reconstructs the class of equivalence of the input via the exploration of the input space by randomly selecting one of the eigenvectors related to the zero eigenvalue. On the opposite, in SiMExp, in order to move from a class of equivalence to another we consider the eigenvectors relative to the nonzero eigenvalues. This requires the slight difference in lines 5 to 7 between Algorithm 1 and Algorithm 2.
1:Set the network $\mathcal{N}$; choose the maximum number of iterations. Choose the input $p_{0}$.
2:for $k=0,1,\dots,K-1$ do
3:    Compute $g_{\mathcal{N}(p_{k})}^{n}$
4:    Compute the pullback metric $g^{0}_{p_{k}}$
5:    Diagonalize $g^{0}_{p_{k}}$ and find the eigenvectors $\{v_{l}\}_{l}$ associated to the zero eigenvalue
6:    Randomly select $\tilde{v}\in\{v_{l}\}_{l}$
7:    $\delta=1/\sqrt{\max(\text{eigenvalues of }g^{0}_{p_{k}})}$
8:    $p_{k+1}\leftarrow p_{k}+\delta\tilde{v}$
9:end for
10:Optionally: store $\{p_{k}\}_{k=0,\dots,K}$ for optimizing future computations
11:Project $p_{k}$ to the nearest feasible region
1:Set the network $\mathcal{N}$; choose the maximum number of iterations. Choose the input $p_{0}$.
2:for $k=0,1,\dots,K-1$ do
3:    Compute $g_{\mathcal{N}(p_{k})}^{n}$
4:    Compute the pullback metric $g^{0}_{p_{k}}$
5:    Diagonalize $g^{0}_{p_{k}}$ and find the eigenvectors $\{w_{l}\}_{l}$ associated to the non-zero eigenvalue
6:    Randomly select $\tilde{w}\in\{w_{l}\}_{l}$
7:    $\delta=2/\sqrt{\max(\text{eigenvalues of }g^{0}_{p_{k}})}$
8:    $p_{k+1}\leftarrow p_{k}+\delta\tilde{w}$
9:end for
10:Optionally: store $\{p_{k}\}_{k=0,\dots,K}$ for optimizing future computations
11:Project $p_{k}$ to the nearest feasible region
Algorithm 2 The Singular Metric Exploration (SiMExp) algorithm.
@@
There are some remarks to point out. From a numerical point of view, the diagonalization of the pullback may lead to have even negative eigenvalues: hence one may use the notion of energy of a curve, related to the pseudolength. The update rule for the new point (line 8) amounts to solve the differential problem via the Euler method: for a reliable solution, we suggest to choose a small step-length $\delta$. On the other hand, if the value of $\delta$ is too small more iterations are needed to move away from the starting point sensibly. Therefore there is a trade-off between the reliability of the solution and the exploration pace. The proof of the well-posedness theorem for Cauchy problems, cf. [18, Theorem 2.1], yields some insights, suggesting to set $\delta$ equal to the inverse of the Lipschitz constant of the map $\mathcal{N}$ – which in practice we can estimate with the inverse of the square root of the largest eigenvalue $\lambda_{M}$ of the pullback metric $g^{0}_{p_{k}}$. This is our default choice for Algorithm 1. We also note that Algorithm 1 is more sensitive to the choice of the parameter $\delta$ compared to Algorithm 2. To build points in the same equivalence class Algorithm 1 needs to follow a null curve closely with as little approximations as possible, namely with a small $\delta$. In contrast Algorithm 2, whose goal is to change the equivalence class from one iteration to the next, does not have the same problem and larger $\delta$ are allowed. Out default choice is therefore to set $\delta=2\lambda_{M}^{-1/2}$ for Algorithm 2. As for the computational complexity of the two algorithms, the most demanding step is the computation of the eigenvalues and eigenvectors, which is $O(n^{3})$, with $n$ the dimension of the square matrix $g^{0}_{p_{k}}$ [20]. Since all the other operations are either $O(n)$ or $O(n^{2})$, we conclude that the complexity of both Algorithms 1 and 2 is $O(n^{3})$.
3.2 Interpretability

### v2

@@ Abstract
This paper introduces a general method for the exploration of equivalence classes in the input space of Transformer models. The proposed approach is based on sound mathematical theory which describes the internal layers of a Transformer architecture as sequential deformations of the input manifold. Using eigendecomposition of the pullback of the distance metric defined on the output space through the Jacobian of the model, we are able to reconstruct equivalence classes in the input space and navigate across them. Our method enables two complementary exploration procedures: the first retrieves input instances that produce the same class probability distribution as the original instance—thus identifying elements within the same equivalence class—while the second discovers instances that yield a different class probability distribution, effectively navigating toward distinct equivalence classes. Finally, we demonstrate how the retrieved instances can be meaningfully interpreted by projecting their embeddings back into a human-readable format. Disclaimer: This paper includes examples of sensitive and very offensive language solely to illustrate the behavior of LLMs exploring the input space.
@@ 1 Introduction
In the literature, the investigation of the input space of Transformers relies on perturbations of input data using heuristic or gradient-based criteria [24, 17, 14], or on the analysis of specific properties of the embedding space [6] via the production of optimal robust explanations and counterfactuals. In this paper, we propose a method for exploring the input space of Transformer models by identifying equivalence classes with respect to their predictions. Our approach is based on sound mathematical theory which describes the internal layers of a Transformer architecture as sequential deformations of the input manifold. Using eigendecomposition of the pullback of the distance metric defined on the output space through the Jacobian of the model, we are able to reconstruct equivalence classes in the input space and navigate in and across them. The equivalence class consists of the counterimage of a particular probability distribution: this means that the elements of an equivalence class, once fed to the Transformer, will provide different realizations from the same probability distribution. Thanks to our approach, we provide two different methods for exploring the embedding space: the Singular Metric Equivalence Class (SiMEC) and the Singular Metric Exploration (SiMExp) algorithms. The first allows for the identification of inputs within the same equivalence class. This means that, given two data inputs $x,x^{\prime}$ identified through the exploration process as belonging to the same equivalence class and a class label $C$, our method guarantees that the Transformers will assign the same probability, modulo numerical approximations, to that label for both inputs: $p(C|x)\approx p(C|x^{\prime})$.
The second method, instead, allows for the exploration of the embedding space starting from an element within an equivalence class and moving towards a different equivalence class. This means that, given two inputs, namely $x,x^{\prime}$, identified in this way and a class label $C$, we guarantee that the Transformer will assign different probabilities to that label for the two inputs, i.e., $p(C|x)\neq p(C|x^{\prime})$. Our experimental data show that this different probability assignment can lead to a change in the most probable class in the Transformers prediction.
In Section˜2, we summarise the mathematical foundations of our approach. In Section˜3, we present our method for the exploration of equivalence classes in the input of the Transformer models. In Section˜4, we empirically investigate the effectiveness and applicability on Transformer models on textual and visual data. In Section˜5, we discuss the relevant literature on embedding space exploration. Finally, in Section˜6, we give our concluding remarks [^1] The code to reproduce our experiments can be found here:https://github.com/alessiomarta/transformers_equivalence_classes..
@@ 2 Preliminaries
We provide in this Section the theoretical foundation of the proposed approach, namely the Geometric Deep Learning framework based on Riemannian Geometry [1, 2].
@@
$M_{0}$$M_{1}$$M_{2}$$\cdots$$M_{n-1}$$M_{n}$$\scriptstyle{\Lambda_{1}}$$\scriptstyle{\Lambda_{2}}$$\scriptstyle{\Lambda_{3}}$$\scriptstyle{\Lambda_{n-1}}$$\scriptstyle{\Lambda_{n}}$ (1)
We call $M_{0}$ the input manifold and $M_{n}$ the output manifold. All the other manifolds of the sequence are called representation manifolds. The maps $\Lambda_{i}$ are the layers of the neural network. We denote by $\mathcal{N}_{(i)}=\Lambda_{n}\circ\dots\circ\Lambda_{i}:M_{i}\to M_{n}$ the mapping from the $i$-th representation layer to the output layer.
@@
A map $\Lambda_{i}:M_{i-1}\rightarrow M_{i}$ is called a smooth layer if it is the restriction to $M_{i-1}$ of a function $\overline{\Lambda}^{(i)}(x):\mathbb{R}^{d_{i-1}}\rightarrow\mathbb{R}^{d_{i}}$ of the form $\overline{\Lambda}^{(i)}_{\alpha}(x)=F^{(i)}_{\alpha}\left(\sum_{\beta}A_{\alpha\beta}^{(i)}x_{\beta}+b^{(i)}_{\alpha}\right)$, for $i=1,\cdots,n$, $x\in\mathbb{R}^{d_{i}}$, $b^{(i)}\in\mathbb{R}^{d_{i}}$ and $A^{(i)}\in\mathbb{R}^{d_{i}\times d_{i-1}}$, with $F^{(i)}:\mathbb{R}^{d_{i}}\rightarrow\mathbb{R}^{d_{i}}$ a diffeomorphism.
@@
Transformers implicitly apply for this framework, since their modules are smooth functions, such as fully connected layers, GeLU and sigmoid activations, thus including also attention layers.
Our aim is to transport the geometric information on the data lying in the output manifold to the input manifold: this allows us to obtain insight on how the network "sees" the input space, how it manipulates it for reaching its final conclusion. For fulfilling this objective, we need several tools from differential geometry. The first key ingredient is the notion of singular Riemannian metric, which has the intuitive meaning of a degenerate scalar product which changes point to point – the starting point for defining a non-Euclidean pseudodistance between points of a manifold.
@@
Without loss of generality, we can assume the following hypotheses on the sequence (1): i) The manifolds $M_{i}$ are open and path-connected sets of dimension $\dim M_{i}=d_{i}$. ii) The maps $\Lambda_{i}$ are $\mathcal{C}^{1}$ submersions. iii) $\Lambda_{i}(M_{i-1})=M_{i}$ for every $i=1,\cdots,n$. iv) The manifold $M_{n}$ is equipped with the structure of Riemannian manifold, with metric $g^{n}$. Definition˜3 naturally leads to the definition of the pseudolength and of energy of a curve.
@@
Let $\gamma:[a,b]\rightarrow\mathbb{R}^{n}$ a curve defined on the interval $a,b]\subset\mathbb{R}$ and $\|v\|_{p}=\sqrt{g_{p}(v,v)}$ the pseudo–norm induced by the pseudo–metric $g_{p}$ at point $p$. Then the pseudolength of $\gamma$ and its energy are defined as
$Pl(\gamma)=\int_{a}^{b}\|\dot{\gamma}(s)\|_{\gamma(s)}ds=\int_{a}^{b}\sqrt{g_{\gamma(s)}(\dot{\gamma}(s),\dot{\gamma}(s))}ds,\qquad E(\gamma)=\int_{a}^{b}\|\dot{\gamma}(s)\|_{\gamma(s)}^{2}ds$ (2)
@@
The second crucial tool is the notion of pullback of a function. Intuitively, given a map $f:M\rightarrow N$ between two manifolds, the pullback operation allows to transfer the geometric information of the output $N$ onto the input $M$ by means of the Jacobian of $f$. More specifically, let $f$ be a function from $\mathbb{R}^{p}$ to $\mathbb{R}^{q}$, and fix the coordinate systems $x=(x_{1},\dots,x_{p})$ and $y=(y_{1},\dots,y_{q})$ on $\mathbb{R}^{p}$ and on $\mathbb{R}^{q}$, respectively. Moreover, we endow $\mathbb{R}^{q}$ with the standard Euclidean metric $g$, whose associated matrix is the identity. The space $\mathbb{R}^{p}$ can then be equipped with the pullback metric $f^{*}g$ whose representation matrix reads as:
@@
Hereafter, we consider in (1) the case $M_{n}=\mathbb{R}^{q}$, equipped with the trivial metric $g^{(n)}=I_{q}$, i.e., the identity. Each manifold $M_{i}$ of the sequence (1) is equipped with a Riemannian singular metric, denoted with $g^{i}$, obtained via the pullback of $\mathcal{N}_{(i)}$. The pseudolength of a curve $\gamma$ on the $i$-th manifold, namely $Pl_{i}(\gamma)$, is computed via the relative metric $g^{i}$ via (2).
@@ 2.1 General results
Let $\gamma:[0,1]\rightarrow M_{i}$ be a piecewise $\mathcal{C}^{1}$ curve. Let $j\in\{i,i+1,\cdots,n\}$ and consider the curve $\gamma_{j}=\Lambda_{j}\circ\cdots\circ\Lambda_{i}\circ\gamma$ on $M_{j}$. Then $Pl_{i}(\gamma)=Pl_{j}(\gamma_{j})$.
@@
The following corollary contains the natural consequences of the previous result; the second point of the claim below is the counterpart of ˜2.
@@
Under the hypothesis of ˜3, one has that ${M_{i}}/{\sim_{i}}={M_{i}}/{\sim_{\mathcal{N}_{i}}}$. Moreover, if two points $p,q\in M_{i}$ are connected by a $\mathcal{C}^{1}$ curve $\gamma:[0,1]\rightarrow M_{i}$ satisfying $\mathcal{N}_{i}(p)=\mathcal{N}_{i}\circ\gamma(s)$ for every $s\in[0,1]$, then they lie in the same class of equivalence.
One is then able to prove that the set $M_{i}/\sim_{i}$ is a smooth manifold:
@@
This last achievement provides practical insights about the projection $\pi_{i}$ on the quotient space, that consists in the building block of the algorithms used for recovering and exploring the foliation in equivalence classes of a neural network.
@@
In [3] it is shown that these results keep to hold true in the case of convolutional layers and residual connections.
@@ 3 Methodology
The results depicted in Section˜2.1 provide powerful tools for investigating how a neural network sees the input space starting from a point $x$. In particular we point out the following remarks: i) If two points $x,y$ belonging to the input manifold $M_{0}$ are such that $x\sim_{0}y$, then $\mathcal{N}(x)=\mathcal{N}(y)$; ii) given a point $z\in M_{n}$, the counterimage $\mathcal{N}^{-1}(z)$ is a smooth manifold, whose connected components are classes of equivalence in $M_{0}$ with respect to $\sim_{0}$, then a necessary condition for two points $x,y\in M_{0}$ to be in the same class of equivalence is that $\mathcal{N}(x)=\mathcal{N}(y)$; iii) any class of equivalence $[x]$, $x\in M_{0}$, is a maximal integral submanifold of $\mathcal{V}M_{0}$. The above observations directly provide with a strategy to build up the equivalence class of an input point $x\in M_{0}$. ˜5 tells us that $\mathcal{V}M_{0}$ is an integrable distribution, with dimension equal to the dimension of the kernel of $g^{0}$: we can hence find $dim(Ker(g^{0}))$ vector fields which are a base for the tangent space of $M_{0}$. This means that we can compute the eigenvalue decomposition of $g^{0}_{x}$ and consider the $L$ linearly independent eigenvectors, namely $\{v_{l}\}_{l=1,\dots,L}$, associated to the null eigenvalue: these eigenvectors depend smoothly on the point, a fact that is not trivial when the matrix associated to the metric depends on several parameters [15]. We can build then all the null curves by randomly selecting one eigenvector $\tilde{v}\in\{v_{l}\}$ and then reconstruct the curve along the direction $\tilde{v}$ from the starting point $x$. From a practical point of view, one is led to solve the Cauchy problem with first order differential equation $\dot{\gamma}=\tilde{v}$ and initial condition $\gamma(0)=x$.
1:Set the network $\mathcal{N}$; choose the number of iterations $K$. Choose the input $x^{(0)}$.
2:for $k=0,1,\dots,K-1$ do
3:  Compute $g_{\mathcal{N}(x^{(k)})}^{n}$
4:  Compute the pullback metric $g^{0}_{x^{(k)}}$
5:  Diagonalize $g^{0}_{x^{(k)}}$ and find the eigenvectors $\{v_{l}\}_{l\in L_{0}}$ associated to the zero eigenvalue $\lambda_{0}$
6:  Randomly select $\tilde{v}\in\{v_{l}\}_{l\in L_{0}}$
7:  $\delta^{(k)}=\eta\sqrt{\min_{j:\lambda_{j}\neq 0}|\lambda_{j}|/\max_{j}|\lambda_{j}|}$
8:  $x^{(k+1)}\leftarrow x^{(k)}+\delta^{(k)}\tilde{v}$
9:end for
10:Project $x^{(k+1)}$ to the feasible region $\mathcal{X}$
1:Set the network $\mathcal{N}$; choose the number of iterations $K$. Choose the input $x^{(0)}$.
2:for $k=0,1,\dots,K-1$ do
3:  Compute $g_{\mathcal{N}(x^{(k)})}^{n}$
4:  Compute the pullback metric $g^{0}_{x^{(k)}}$
5:  Diagonalize $g^{0}_{x^{(k)}}$ and find the eigenvectors $\{w_{l}\}_{l\in L_{+}}$ associated to eigenvalues $\lambda_{l}\neq 0$
6:  Randomly select $\tilde{w}\in\{w_{l}\}_{l\in L_{+}}$
7:  $\delta^{(k)}=\eta\sqrt{\min_{j:\lambda_{j}\neq 0}|\lambda_{j}|/\max_{j}|\lambda_{j}|}$
8:  $x^{(k+1)}\leftarrow x^{(k)}+\delta^{(k)}\tilde{w}$
9:end for
10:Project $x^{(k+1)}$ to the feasible region $\mathcal{X}$
Algorithm 2 The Singular Metric Exploration (SiMExp) algorithm.
@@ 3.1 Input Space Exploration
This entire procedure is coded in the Singular Metric Equivalence Class (SiMEC) and Singular Metric Exploration (SiMExp) algorithms, whose general schemes are depicted in Algorithms˜1 and 2. SiMEC reconstructs the class of equivalence of the input via exploration of the input space by randomly selecting one of the eigenvectors related to the zero eigenvalue. On the opposite, in SiMExp, in order to move from a class of equivalence to another we consider the eigenvectors relative to the nonzero eigenvalues. This requires the slight difference in lines 5-6 between Algorithm˜1 and Algorithm˜2.
@@
There are some remarks to point out. From a numerical point of view, the diagonalization of the pullback may lead to have even negative eigenvalues: hence one may use the notion of energy of a curve, related to the pseudolength. If the values $\delta^{(k)}$ are too small more iterations are needed to move away from the starting point sensibly. Therefore there is a trade-off between the reliability of the solution and the exploration pace. Relying on the theory of dynamical systems, we can in practice estimate $\delta^{(k)}$ at each iteration with the inverse of the square root of the condition number $\Gamma=\max_{j}|\lambda_{j}|/\min_{j:\lambda_{j}\neq 0}|\lambda_{j}|$ of the pullback metric $g^{0}_{x^{(k)}}$, as in a locally-linearized dynamical system. This is our default choice for both algorithms. We multiply the default value $\delta$ by a multiplier $\eta$ in order to explore the sensitivity of Algorithm˜1 and Algorithm˜2 to variations of step length, expecting Algorithm˜1 to be more sensitive compared to Algorithm˜2. Indeed, to build points in the same equivalence class Algorithm˜1 needs to follow a null curve closely with as little approximation as possible. In contrast Algorithm˜2, whose goal is to change the equivalence class, does not have the same problem and larger $\delta$ are allowed.
[…]
As for the computational complexity of the two algorithms, the most demanding step is the computation of the eigenvalues and eigenvectors, which is $O(n^{3})$, with $n$ the dimension of the square matrix $g^{0}_{x^{(k)}}$ [20]. Since all the other operations are either $O(n)$ or $O(n^{2})$, we conclude that the complexity of both Algorithms˜1 and 2 is $O(n^{3})$.
3.2 Interpretation